Create a real life situation that fits into the equation (x+3)(x-5)=0 and express the situation as the same equation.
Answer by Alan3354(69443)

Create a real life situation that fits into the equation (x+3)(x-5)=0 and express the situation as the same equation.
------------------
I just did a problem like that. Change the 150 to 15 and it fits.
Steve traveled 150 miles at a certain speed. Had he gone 20mph faster, the trip would have taken 2 hours less. Find the speed of his vehicle.
----------------
t = 150/r t=time, r=rate
t-2 = 150/(r+20)
Sub for t in 2nd eqn
(150/r)-2 = 150/(r+20)
Multiply by (r+20)
(150(r+20)/r)-2(r+20) = 150
Multiply by r
150(r+20) - 2r(r+20) = 150r
(r+50)*(r-30) = 0
r = 30, r = -50 *********(Ignore the -50)
r = 30 mph